### v1

# Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation

Abstract
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements in integrating different modalities, excelling in complex understanding and generation tasks. Despite their success, MLLMs remain vulnerable to conversational adversarial inputs, particularly negation arguments. This paper systematically evaluates state-of-the-art MLLMs across diverse benchmarks, revealing significant performance drops when negation arguments are introduced to initially correct responses. We show critical vulnerabilities in the reasoning and alignment mechanisms of these models. Proprietary models such as GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA. However, all evaluated MLLMs struggle to maintain logical consistency under negation arguments during conversation. This paper aims to offer valuable insights for improving the robustness of MLLMs against adversarial inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
Figure 1: Examples demonstrate that GPT-4o initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments. GPT-4o shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting".
1 Introduction
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt . These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning. Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning. Additionally, open-source models like DeepSeek liu2024deepseek are closing the performance gap with the proprietary models, reflecting the rapid and competitive evolution of MLLMs. These advancements not only highlight the growing capabilities of MLLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation. This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning . As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples. However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly. Moreover, it even generates hallucinated explanations to justify these erroneous responses. Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms. These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial. Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
Figure 2: Performance comparison of selected Multimodal Large Language Models (MLLMs) across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista )-showing average accuracy before (initial answers) and after the introduction of negation arguments. Note that GPT-4o (accessed from 20-Jan-2025 to 22-Jan-2025) and Claude-3.5-Sonnet (accessed from 20-Jan-2025 to 23-Jan-2025) results are averaged over six benchmarks, excluding POPE and MMBench. Results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025. More detail results can be found in Table 1.
To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across multiple major benchmarks and state-of-the-art models. The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation. The selected proprietary models, such as GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy. Through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
Our observations reveal critical insights into the behavior of MLLMs. Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, suggesting potential overfitting or alignment issues during training. On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments. These findings emphasize the urgent need for enhanced adversarial training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
2 Evaluating Multimodal Large Language Model via Negation
2.1 Experimental Setup
Multimodal Large Language Models. To establish a comprehensive evaluation of the negation challenges posed to Multimodal Large Language Models (MLLMs), we assess a range of state-of-the-art models representing both proprietary and open-source systems. These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 . These models represent the most powerful multimodal AI capabilities.
Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B . These open-source models showcase the rapid progress in publicly available multimodal AI systems.
Benchmarks. To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
Evaluation Pipeline. To evaluate the robustness of MLLMs, we employ a structured evaluation pipeline designed to measure the impact of negation on model performance. The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded. Second, if the model’s initial answer aligns with the correct answer from the dataset, we introduce a negation argument challenging the validity of the original answer. The model’s response to the negation argument is recorded to assess whether it revises its original correct answer. If the model’s initial answer is incorrect, no negation argument is introduced, and the response remains unchanged. Third, we compute the accuracy of the model both before and after the introduction of negation arguments to quantify the drop in performance caused by adversarial negation. This evaluation pipeline enables a detailed analysis of how well MLLMs can maintain consistent reasoning under adversarial conversational challenges, shedding light on their robustness and potential vulnerabilities.
Negation Argument Generation. We systematically generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions. Negation arguments are generated by rewriting the original question into an opposite statement based on the answer. If the answer is “Yes.", the question is rewritten as a negative statement and vice versa. For example, given the question “Does this image describe a place of banquet hall?" and the answer “Yes.", the negation argument is “This image does not describe a place of banquet hall." (2) Multiple-Choice Questions. For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer. Given the question and options, “What organism does this diagram portray? Options: A. ladybug B. wolf C. human D. plant", as well as the answer “D". The negation argument could be “The diagram portrays a ladybug." (option A). (3) Free-Form Questions. For free-form answers, negation arguments are crafted to contradict the provided factual answer. For instance, give the question “How many food items are shown in the bar graph? " and answer “14", the negation argument is “The number of food items shown in the bar graph is not 14."
Post-processing. To ensure consistency in evaluation, we employ a post-processing step to handle variations in the responses generated by MLLMs. Since the model outputs may not always align exactly with the expected format, we utilized Llama3-8B-Instruct and Qwen-14B-Chat to refine the responses. These models are tasked with interpreting the question, the expected answer, and the MLLM-generated response, normalizing the output to align with the desired format.
2.2 Results
Model Negation Dataset average
MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
Open-Source
before 79.70 31.77 12.37 59.52 33.10 51.72 86.24 68.79 52.90
[…]
before 86.02 50.37 27.05 78.37 60.00 75.32 87.94 82.26 68.42
[…]
before 70.01 42.98 12.14 67.81 34.90 60.72 88.96 79.44 57.12
[…]
before 91.70 60.34 33.64 83.42 67.10 80.60 87.84 79.81 73.06
[…]
Proprietary
before 82.43 57.39 31.98 71.99 48.70 68.37 80.27 78.84 65.00
[…]
before 68.45 60.47 43.12 78.30 59.00 72.28 63.60
[…]
before 84.46 60.34 51.56 70.63 60.30 78.84 67.69
[…]
Table 1: Performance comparison of selected Multimodal Large Language Models (MLLMs) across various benchmarks before (i.e., initial answers) and after the introduction of negation arguments. The performance drop is highlighted in red.
Performance Comparison. Table 1 lists the results across all evaluated benchmarks, comparing the accuracy of selected MLLMs before (i.e., initial answers) and after the introduction of negation arguments. Overall, all the MLLMs exhibit significant performance declines even though the accuracy of accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts. The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60 (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
Open-Source vs. Proprietary Models. Proprietary models generally showed better resilience. For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.61%, maintaining 44.39% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.15% drop and only manages to obtain the 26.91% accuracy on average. This highlights the need for improved adversarial training in open-source MLLMs to deal with the negation arguments in conversation. In addition, proprietary models like GPT-4o do not consistently outperform open-source models across all datasets. On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown." For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as "unknown." Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
Comparison between Larger and Smaller Models. Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.06%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%. However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.15% decline compared to Qwen2-VL-7B-Instruct’s 32.15% drop. This suggests that larger models may be more prone to overfitting or misalignment in adversarial scenarios like negation.
Qualitative Analysis. Figure 3 illustrates a series of examples showcasing how MLLMs respond to negation arguments in different datasets. In each example, the models initially provide correct responses. However, when users introduce negation arguments, many models revise their answers incorrectly. For instance, GPT-4o initially identifies the color representing New Zealand correctly but revises its answer after the negation argument. In addition, the models often generate detailed but fabricated explanations to justify their revised answers. For example, in the fourth example in MMMU dataset, Gemini-1.5-flash generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
Figure 3: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct. Correct responses are highlighted in green, while incorrect responses are marked in red.
Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
before negation 31.77 42.98 50.37 57.39
[…]
Table 2: Performance comparison of selected Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset. The performance drop compared to before negation is indicated in red.
Effect of Negation Argument. To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options. This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 2, all the MLLMs show higher accuracy for the second strategy after negation. The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments. Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments. Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods. Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
3 Related Work
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities. The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets. Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning. Open-source counterparts, such as LLaVA li2024llavanext-strong , Qwen2-VL Qwen2VL-7B and Deepseek liu2024deepseek , have democratized access to advanced multimodal technologies, fostering innovation within the research community. The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions. Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct. Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
Negation Understanding. Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development . In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert . To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding . This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation. These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts. Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can . Very recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments. Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
In multimodal learning, negation has received comparatively less attention than in NLP. Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision . For instance, very recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance. In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
4 Conclusion
This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct. While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms. Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs. This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial conditions. We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.

### v2

@@
 This paper systematically evaluates state-of-the-art MLLMs across diverse benchmarks, revealing significant performance drops when negation arguments are introduced to initially correct responses.
-We show critical vulnerabilities in the reasoning and alignment mechanisms of these models.
-Proprietary models such as GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA.
+Notably, we introduce the first benchmark GaslightingBench, specifically designed to evaluate the vulnerability of MLLMs to negation arguments.
+GaslightingBench consists of multiple-choice questions curated from existing datasets, along with generated negation prompts across 20 diverse categories.
+Throughout extensive evaluation, we find that proprietary models such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA.
 However, all evaluated MLLMs struggle to maintain logical consistency under negation arguments during conversation.
-This paper aims to offer valuable insights for improving the robustness of MLLMs against adversarial inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
+Our findings provide critical insights for improving the robustness of MLLMs against negation inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
+1 Introduction
 Figure 1: Examples demonstrate that GPT-4o initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments.
 GPT-4o shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting".
-1 Introduction
-Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt .
-These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning.
-Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning.
-Additionally, open-source models like DeepSeek liu2024deepseek are closing the performance gap with the proprietary models, reflecting the rapid and competitive evolution of MLLMs.
-These advancements not only highlight the growing capabilities of MLLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
+Note: "Gaslighting is the manipulation of someone into questioning their own perception of reality." - Wikipedia.
+Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs [32, 19, 3, 10].
+These models leverage cutting-edge advancements in Large Language Models (LLM) [36] and multimodal learning [25, 14, 17], enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning.
+Recent breakthroughs, such as the introduction of proprietary models GPT-4o [10] and Claude-3.5-Sonnet [2], have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning.
+Additionally, open-source models like DeepSeek-V3 and DeepSeek-R1 [16, 7] are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs.
+These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
 Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation.
-This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning .
+This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers [29, 37].
 As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples.
@@
 These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial.
-Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
-Figure 2: Performance comparison of selected Multimodal Large Language Models (MLLMs) across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista )-showing average accuracy before (initial answers) and after the introduction of negation arguments.
-Note that GPT-4o (accessed from 20-Jan-2025 to 22-Jan-2025) and Claude-3.5-Sonnet (accessed from 20-Jan-2025 to 23-Jan-2025) results are averaged over six benchmarks, excluding POPE and MMBench.
-Results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025.
-More detail results can be found in Table 1.
-To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across multiple major benchmarks and state-of-the-art models.
-The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation.
-The selected proprietary models, such as GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy.
-Through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
+Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models [9, 20].
+Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME [6], MMMU [33], MMMUPro [34], MMBench [21], PoPE [15], ChartQA [23], AI2Diagram [12] and MathVista [22]).
+The results highlight the substantial accuracy drop across all models when negation is introduced.
+The access dates for GPT-4o and Claude-3.5-Sonnet are from 3-March-2025 to 7-March-2025, while results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025.
+More detailed results are available in Table 1.
+To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across eight multimodal benchmarks and state-of-the-art models.
+The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU [33] and MMBench [21], chart dataset such as ChartQA [23], to math dataset MathVista [22], providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation.
+The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy.
+As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
+As existing benchmarks primarily evaluate factual accuracy and multimodal reasoning, they fail to systematically assess MLLMs’ susceptibility to negation arguments.
+To address this gap, we introduce GaslightingBench, the first multimodal benchmark designed to evaluate models’ ability to resist misleading negation arguments while maintaining logical consistency.
+We curate representative multiple-choice questions (MCQs) from established datasets and generated corresponding negation prompts, resulting in a collection spanning 20 categories and 1,287 samples.
 Our observations reveal critical insights into the behavior of MLLMs.
@@
 On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments.
-These findings emphasize the urgent need for enhanced adversarial training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
+Furthermore, different categories demonstrate varying levels of vulnerability to negation, with “Social Relation" suffering the most significant performance declines, whereas “Geography" experiences comparatively smaller yet still notable drops in our GaslightingBench.
+These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
+2 Related Work
+Multimodal Large Language Models.
+Multimodal Large Language Models (MLLMs) [32] represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities.
+The introduction of contrastive learning techniques in vision-language models such as CLIP [25] enables cross-modal understanding through training on extensive datasets.
+Building upon these foundations, modern MLLMs like GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27] have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning.
+Open-source counterparts, such as LLaVA [13] and Qwen2-VL [31], have democratized access to advanced multimodal technologies, fostering innovation within the research community.
+The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions.
+Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct.
+Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
+Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to negation.
+The pipeline consists of three key stages: (1) Inputs and Initial Answers: MLLMs receive a variety of question formats as input, including Yes/No, Multiple-Choice, and Free-Form, and their initial answers are recorded.
+(2) Negation Generation: if the model’s initial response is correct, a negation argument is introduced to challenge its answer.
+Different negation strategies are applied based on the question type.
+(3) Post-Negation Evaluation: the model’s response after negation is analyzed to determine if it maintains consistency or is misled into revising its answer.
+Post-processing is applied to normalize responses for accurate comparison.
+Negation Understanding.
+Negation, defined as the contradiction or denial of something, is a fundamental aspect of language [4, 24].
+In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions [11, 5].
+To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets [8].
+This line of research has advanced further with studies like [28], which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation.
+These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts.
+Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding [29].
+Very recent work [37] addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments.
+Building on the insights of [29, 37], this paper extends the investigation into the impact of opposing arguments in MLLMs.
+In multimodal learning, negation has received comparatively less attention than in NLP.
+Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) [35, 26, 30, 1].
+For instance, very recent work [1] highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance.
+In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
 3 Evaluating Multimodal Large Language Model via Negation
 3.1 Experimental Setup
+Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples.
+Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to negation arguments.
+Figure 5: Examples from different categories in the GaslightingBench.
+The highlighted green option is the correct answer, while a randomly selected incorrect option is used to generate the negation argument.
 Multimodal Large Language Models.
@@
 These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
-Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 .
+Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27].
 These models represent the most powerful multimodal AI capabilities.
-Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B .
+Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B [18], Qwen2-VL-7B-Instruct [31], LLaVA-NeXT-8B [13] and Qwen2-VL-72B-Instruct [31].
 These open-source models showcase the rapid progress in publicly available multimodal AI systems.
 Benchmarks.
-To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
-Evaluation Pipeline.
-To evaluate the robustness of MLLMs, we employ a structured evaluation pipeline designed to measure the impact of negation on model performance.
+To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME [6], MMMU [33], MMMUPro [34], MMBench [21] and PoPE [15], Chart dataset ChartQA [23], diagram dataset AI2Diagram [12] as well as Math dataset MathVista [22].
+3.2 Evaluation Pipeline
+As shown in Figure 3, we introduce a structured evaluation pipeline to systematically assess MLLMs’ vulnerability to negation by measuring their performance before and after exposure to negation arguments.
 The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded.
@@
 Third, we compute the accuracy of the model both before and after the introduction of negation arguments to quantify the drop in performance caused by adversarial negation.
-This evaluation pipeline enables a detailed analysis of how well MLLMs can maintain consistent reasoning under adversarial conversational challenges, shedding light on their robustness and potential vulnerabilities.
-Negation Argument Generation.
-We systematically generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions.
+Negation Generation.
+As illustrated in Figure 3, we generate negation arguments tailored to different question formats using Llama3-8B-Instruct: (1) Yes/No Questions.
 Negation arguments are generated by rewriting the original question into an opposite statement based on the answer.
 If the answer is “Yes.", the question is rewritten as a negative statement and vice versa.
-For example, given the question “Does this image describe a place of banquet hall?" and the answer “Yes.", the negation argument is “This image does not describe a place of banquet hall." (2) Multiple-Choice Questions.
+For example, given the question “Is this photo of Grimspound?" and the answer “Yes.", the negation argument is “This it not a photo of Grimspound." (2) Multiple-Choice Questions.
 For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer.
-Given the question and options, “What organism does this diagram portray?
+Given the question and options, “Which direction is the cat facing?
 Options: A.
-ladybug B.
-wolf C.
-human D.
-plant", as well as the answer “D".
-The negation argument could be “The diagram portrays a ladybug." (option A).
+right B.
+upward C.
+left D.
+the camera", as well as the answer “D".
+The negation argument could be “The cat is facing right." (option A).
 (3) Free-Form Questions.
 For free-form answers, negation arguments are crafted to contradict the provided factual answer.
-For instance, give the question “How many food items are shown in the bar graph?
-" and answer “14", the negation argument is “The number of food items shown in the bar graph is not 14."
+For instance, give the question “What is the maximum value of y?
+" and answer “5", the negation argument is “The maximum value of y is not 5."
+Model Negation Dataset average
+MME [6] MMMU [33] MMMUPro [34] AI2Diagram [12] MathVista [22] ChartQA [23] PoPE [15] MMBench [21]
+Open-Source
+before 79.70 31.77 12.37 60.40 33.10 51.72 86.24 68.79 53.01
+[…]
+before 86.02 50.37 27.05 79.83 60.00 75.32 87.94 82.26 68.42
+[…]
+before 70.01 42.98 12.14 68.56 34.90 60.72 88.96 79.44 57.21
+[…]
+before 91.70 60.34 33.64 85.30 67.10 80.60 87.84 79.81 73.29
+[…]
+Proprietary
+before 82.43 57.39 31.98 73.52 48.70 68.37 80.27 78.84 65.19
+[…]
+before 69.17 62.07 36.42 75.84 54.60 72.68 85.46 77.69 66.74
+[…]
+before 86.60 67.73 37.40 72.51 57.50 77.40 81.71 83.07 70.49
+[…]
+Table 1: Performance comparison of Multimodal Large Language Models (MLLMs) across various benchmarks before (i.e., initial answers) and after the introduction of negation arguments.
+The performance drop is highlighted in red.
+Model LLaVA-1.6-7B [18] Qwen2-VL-7B-Instruct [31] LLaVA-NeXT-8B [13] Qwen2-VL-72B-Instruct [31] Gemini-1.5-flash [27] GPT-4o [10] Claude-3.5-Sonnet [2]
+before negation 59.13 76.85 71.33 77.08 74.13 69.15 78.17
+[…]
+Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after negation. The performance drop is highlighted in red.
+Model LLaVA-1.6-7B [18] LLaVA-NeXT-8B [13] Qwen2-VL-7B-Instruct [31] Gemini-1.5-flash [27]
+before negation 31.77 42.98 50.37 57.39
+[…]
+[…]
+Table 3: Performance comparison of Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset.
+The performance drop compared to before negation is indicated in red.
 Post-processing.
@@
 These models are tasked with interpreting the question, the expected answer, and the MLLM-generated response, normalizing the output to align with the desired format.
-2.2 Results
-Model Negation Dataset average
-MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
-Open-Source
-before 79.70 31.77 12.37 59.52 33.10 51.72 86.24 68.79 52.90
-[…]
-before 86.02 50.37 27.05 78.37 60.00 75.32 87.94 82.26 68.42
-[…]
-before 70.01 42.98 12.14 67.81 34.90 60.72 88.96 79.44 57.12
-[…]
-before 91.70 60.34 33.64 83.42 67.10 80.60 87.84 79.81 73.06
-[…]
-Proprietary
-before 82.43 57.39 31.98 71.99 48.70 68.37 80.27 78.84 65.00
-[…]
-before 68.45 60.47 43.12 78.30 59.00 72.28 63.60
-[…]
-before 84.46 60.34 51.56 70.63 60.30 78.84 67.69
-[…]
-Table 1: Performance comparison of selected Multimodal Large Language Models (MLLMs) across various benchmarks before (i.e., initial answers) and after the introduction of negation arguments.
-The performance drop is highlighted in red.
-Performance Comparison.
+3.3 Gaslighting Benchmark Construction
+To comprehensively evaluate the impact of negation on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments.
+GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions.
+Our benchmark is constructed through a two-step process.
+We first select representative questions and images.
+MCQs are extracted from existing datasets following three key principles: (1) We carefully review the categories across all the existing benchmarks and select a balanced representation of general categories.
+(2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories.
+For example, in MMBench,“Attibute_comparison" and “Attribute_recognition" are combined under “Attribute" to ensure broader coverage.
+(3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness.
+For instance, our “Math" category consists of samples from both AI4MATH and MMMU-Math.
+Second, we employ the same approach detailed in section 3.2 to generate negation arguments for all the selected questions.
+Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting.
+Figure 5 shows a few examples from different categories in the GaslightingBench.
+4 Result Analysis
+Figure 6: Accuracy drop across different categories in GaslightingBench.
+4.1 Performance Comparison
 Table 1 lists the results across all evaluated benchmarks, comparing the accuracy of selected MLLMs before (i.e., initial answers) and after the introduction of negation arguments.
-Overall, all the MLLMs exhibit significant performance declines even though the accuracy of accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts.
-The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60 (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
+Overall, all the MLLMs exhibit significant performance declines even though the accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts.
+The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60% (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
 Open-Source vs.
@@
 Proprietary models generally showed better resilience.
-For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.61%, maintaining 44.39% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.15% drop and only manages to obtain the 26.91% accuracy on average.
+For instance, Gemini-1.5-flash has the lowest average accuracy drop 20.53%, maintaining 44.66% accuracy after negation, whereas open-source models like Qwen2-VL-72B-Instruct suffered dramatic declines, with a 46.02% drop and only manages to obtain 27.27% accuracy on average.
 This highlights the need for improved adversarial training in open-source MLLMs to deal with the negation arguments in conversation.
@@
 Comparison between Larger and Smaller Models.
-Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.06%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%.
-However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.15% decline compared to Qwen2-VL-7B-Instruct’s 32.15% drop.
+Larger models like Qwen2-VL-72B-Instruct performed significantly better initially before negation, achieving an average accuracy of 73.29%, compared to smaller models like Qwen2-VL-7B-Instruct with 68.42%.
+However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.02% decline compared to Qwen2-VL-7B-Instruct’s 32.17% drop.
 This suggests that larger models may be more prone to overfitting or misalignment in adversarial scenarios like negation.
+Figure 7: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct.
+Correct responses are highlighted in green, while incorrect responses are marked in red.
+Result in GaslightingBench.
+Table 2 presents the evaluation results of MLLMs on GaslightingBench, highlighting their accuracy before and after the introduction of negation arguments.
+Similarly, across all models, significant performance degradation is observed, reaffirming the susceptibility of MLLMs to negation.
+As shown in Figure 6, different categories show varying degrees of susceptibility to negation arguments.
+Notably, “Social Relation" and “Image Emotion" experience the most severe drops, with accuracy declining by over 50%, indicating their high sensitivity to negation arguments.
+In contrast, “Geography" and “Chemistry" demonstrate comparatively lower declines but still suffer more than a 25% accuracy reduction, emphasizing that no category remains unaffected.
+Similar observations also appear in other existing datasets, for instance, MMMU dataset, which contains 30 subjects such as “accounting" and “Art”, the largest performance drop is observed in “Art Theory", while “Music" experiences the smallest decline.
+Similarly, in the MME dataset, “Counting" shows the most significant drop, whereas “Numerical Calculation" is the least affected.
+These results highlight the uneven but widespread impact of negation arguments across different domains.
 Qualitative Analysis.
-Figure 3 illustrates a series of examples showcasing how MLLMs respond to negation arguments in different datasets.
+Figure 7 illustrates examples showcasing how MLLMs respond to negation arguments in different datasets.
 In each example, the models initially provide correct responses.
 However, when users introduce negation arguments, many models revise their answers incorrectly.
-For instance, GPT-4o initially identifies the color representing New Zealand correctly but revises its answer after the negation argument.
+For instance, in the first example in ChartQA dataset, GPT-4o initially identifies the color representing New Zealand correctly but revises its answer after the negation argument.
 In addition, the models often generate detailed but fabricated explanations to justify their revised answers.
-For example, in the fourth example in MMMU dataset, Gemini-1.5-flash generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
-Figure 3: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct.
-Correct responses are highlighted in green, while incorrect responses are marked in red.
-Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
-before negation 31.77 42.98 50.37 57.39
-[…]
-[…]
-Table 2: Performance comparison of selected Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset.
-The performance drop compared to before negation is indicated in red.
-Effect of Negation Argument.
+For example, in the second example in MMMU dataset, LLaVA-1.6-7B generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
+Figure 8: Example shows that different options in negation arguments (e.g., “ten," “one," and “eight") affect MLLMs’ ability to maintain their initial correct response.
+4.2 Discussion: Effect of Negation Argument
 To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options.
@@
 Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments.
+Furthermore, as shown in Figure 8, the choice of negation arguments significantly affects MLLMs’ performance.
+For instance, when the negation argument includes “ten" or “eight," only two models maintain the correct response of “four," whereas using “one" as the negation argument allows six models to retain the correct answer.
+This highlights the varying sensitivity of MLLMs to different negation formulations.
 Additionally, we acknowledge that negation arguments can take numerous other forms, such as employing alternative phrasing methods.
 Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
-3 Related Work
-Multimodal Large Language Models.
-Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities.
-The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets.
-Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning.
-Open-source counterparts, such as LLaVA li2024llavanext-strong , Qwen2-VL Qwen2VL-7B and Deepseek liu2024deepseek , have democratized access to advanced multimodal technologies, fostering innovation within the research community.
-The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions.
-Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct.
-Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
-Negation Understanding.
-Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development .
-In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert .
-To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding .
-This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation.
-These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts.
-Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can .
-Very recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments.
-Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
-In multimodal learning, negation has received comparatively less attention than in NLP.
-Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision .
-For instance, very recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance.
-In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
 5 Conclusion
@@
 Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs.
-This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial conditions.
+This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions.
 We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.

### v3

@@
 Note: "Gaslighting is the manipulation of someone into questioning their own perception of reality." - Wikipedia.
-Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs [32, 19, 3, 10].
-These models leverage cutting-edge advancements in Large Language Models (LLM) [36] and multimodal learning [25, 14, 17], enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning.
-Recent breakthroughs, such as the introduction of proprietary models GPT-4o [10] and Claude-3.5-Sonnet [2], have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning.
-Additionally, open-source models like DeepSeek-V3 and DeepSeek-R1 [16, 7] are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs.
+Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt .
+These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning.
+Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning.
+Additionally, open-source models like DeepSeek-V3, DeepSeek-R1 liu2024deepseek ; guo2025deepseek and Qwen series bai2023qwen ; qwen2.5 are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs.
 These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
 Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation.
-This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers [29, 37].
+This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning .
 As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples.
@@
 Moreover, it even generates hallucinated explanations to justify these erroneous responses.
-Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms.
+Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms sharmatowards .
 These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial.
-Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models [9, 20].
-Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME [6], MMMU [33], MMMUPro [34], MMBench [21], PoPE [15], ChartQA [23], AI2Diagram [12] and MathVista [22]).
+Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
+Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista ).
 The results highlight the substantial accuracy drop across all models when negation is introduced.
@@
 More detailed results are available in Table 1.
-To comprehensively understand the limitations of MLLMs in handling negation, we conduct extensive evaluations across eight multimodal benchmarks and state-of-the-art models.
-The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU [33] and MMBench [21], chart dataset such as ChartQA [23], to math dataset MathVista [22], providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation.
+To comprehensively understand the limitations of MLLMs in handling negation, this paper presents the first systematic study by conducting extensive evaluations across eight multimodal benchmarks and state-of-the-art models.
+The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation.
 The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy.
@@
 Our observations reveal critical insights into the behavior of MLLMs.
-Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, suggesting potential overfitting or alignment issues during training.
+Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, indicating a lack of robustness in handling negation inputs.
 On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments.
@@
 These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
-2 Related Work
+2 Evaluating Multimodal Large Language Model via Negation
+2.1 Experimental Setup
 Multimodal Large Language Models.
-Multimodal Large Language Models (MLLMs) [32] represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities.
-The introduction of contrastive learning techniques in vision-language models such as CLIP [25] enables cross-modal understanding through training on extensive datasets.
-Building upon these foundations, modern MLLMs like GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27] have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning.
-Open-source counterparts, such as LLaVA [13] and Qwen2-VL [31], have democratized access to advanced multimodal technologies, fostering innovation within the research community.
-The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions.
-Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct.
-Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
+To establish a comprehensive evaluation of the negation challenges posed to Multimodal Large Language Models (MLLMs), we assess a range of state-of-the-art models representing both proprietary and open-source systems.
+These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
+Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 .
+These models represent the most powerful multimodal AI capabilities.
+All evaluations for proprietary models are performed via their publicly available APIs.
+Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B .
+These open-source models showcase the rapid progress in publicly available multimodal AI systems.
+All evaluations for these models are conducted using a single NVIDIA H100 GPU.
+Benchmarks.
+To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
 Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to negation.
@@
 Post-processing is applied to normalize responses for accurate comparison.
-Negation Understanding.
-Negation, defined as the contradiction or denial of something, is a fundamental aspect of language [4, 24].
-In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions [11, 5].
-To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets [8].
-This line of research has advanced further with studies like [28], which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation.
-These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts.
-Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding [29].
-Very recent work [37] addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments.
-Building on the insights of [29, 37], this paper extends the investigation into the impact of opposing arguments in MLLMs.
-In multimodal learning, negation has received comparatively less attention than in NLP.
-Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) [35, 26, 30, 1].
-For instance, very recent work [1] highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance.
-In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
-3 Evaluating Multimodal Large Language Model via Negation
-3.1 Experimental Setup
-Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples.
-Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to negation arguments.
-Figure 5: Examples from different categories in the GaslightingBench.
-The highlighted green option is the correct answer, while a randomly selected incorrect option is used to generate the negation argument.
-Multimodal Large Language Models.
-To establish a comprehensive evaluation of the negation challenges posed to Multimodal Large Language Models (MLLMs), we assess a range of state-of-the-art models representing both proprietary and open-source systems.
-These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
-Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o [10], Claude-3.5-Sonnet [2], and Gemini-1.5-flash [27].
-These models represent the most powerful multimodal AI capabilities.
-Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B [18], Qwen2-VL-7B-Instruct [31], LLaVA-NeXT-8B [13] and Qwen2-VL-72B-Instruct [31].
-These open-source models showcase the rapid progress in publicly available multimodal AI systems.
-Benchmarks.
-To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME [6], MMMU [33], MMMUPro [34], MMBench [21] and PoPE [15], Chart dataset ChartQA [23], diagram dataset AI2Diagram [12] as well as Math dataset MathVista [22].
 2.2 Evaluation Pipeline
@@
 " and answer “5", the negation argument is “The maximum value of y is not 5."
+Post-processing.
+To ensure consistency in evaluation, we employ a post-processing step to handle variations in the responses generated by MLLMs.
+Since the model outputs may not always align exactly with the expected format, we utilized Llama3-8B-Instruct and Qwen-14B-Chat to refine the responses.
+These models are tasked with interpreting the question, the expected answer, and the MLLM-generated response, normalizing the output to align with the desired format.
+Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples.
+Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to negation arguments.
+Figure 5: Examples from different categories in the GaslightingBench. The green-highlighted option is correct, while a randomly chosen incorrect option is used to generate the negation argument.
+2.3 Gaslighting Benchmark Construction
+To comprehensively evaluate the impact of negation on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments.
+GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions.
+Our benchmark is constructed through a two-step process.
+We first select representative questions and images.
+MCQs are extracted from existing datasets following three key principles: (1) We carefully review the categories across all the existing benchmarks and select a balanced representation of general categories.
+(2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories.
+For example, in MMBench,“Attibute_comparison" and “Attribute_recognition" are combined under “Attribute" to ensure broader coverage.
+(3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness.
+For instance, our “Math" category consists of samples from both AI4MATH and MMMU-Math.
+Second, we employ the same approach detailed in section 2.2 to generate negation arguments for all the selected questions.
+Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting.
+Figure 5 shows a few examples from different categories in the GaslightingBench.
+3 Result Analysis
 Model Negation Dataset average
-MME [6] MMMU [33] MMMUPro [34] AI2Diagram [12] MathVista [22] ChartQA [23] PoPE [15] MMBench [21]
+MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
 Open-Source
@@
 The performance drop is highlighted in red.
-Model LLaVA-1.6-7B [18] Qwen2-VL-7B-Instruct [31] LLaVA-NeXT-8B [13] Qwen2-VL-72B-Instruct [31] Gemini-1.5-flash [27] GPT-4o [10] Claude-3.5-Sonnet [2]
+Model LLaVA-1.6-7B liu2024llavanext-llava1.6 Qwen2-VL-7B-Instruct Qwen2VL-7B LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-72B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5 GPT-4o hurst2024gpt Claude-3.5-Sonnet anthropic2024claude
 before negation 59.13 76.85 71.33 77.08 74.13 69.15 78.17
@@
 Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after negation. The performance drop is highlighted in red.
-Model LLaVA-1.6-7B [18] LLaVA-NeXT-8B [13] Qwen2-VL-7B-Instruct [31] Gemini-1.5-flash [27]
+Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
 before negation 31.77 42.98 50.37 57.39
@@
 The performance drop compared to before negation is indicated in red.
-Post-processing.
-To ensure consistency in evaluation, we employ a post-processing step to handle variations in the responses generated by MLLMs.
-Since the model outputs may not always align exactly with the expected format, we utilized Llama3-8B-Instruct and Qwen-14B-Chat to refine the responses.
-These models are tasked with interpreting the question, the expected answer, and the MLLM-generated response, normalizing the output to align with the desired format.
-3.3 Gaslighting Benchmark Construction
-To comprehensively evaluate the impact of negation on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments.
-GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions.
-Our benchmark is constructed through a two-step process.
-We first select representative questions and images.
-MCQs are extracted from existing datasets following three key principles: (1) We carefully review the categories across all the existing benchmarks and select a balanced representation of general categories.
-(2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories.
-For example, in MMBench,“Attibute_comparison" and “Attribute_recognition" are combined under “Attribute" to ensure broader coverage.
-(3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness.
-For instance, our “Math" category consists of samples from both AI4MATH and MMMU-Math.
-Second, we employ the same approach detailed in section 3.2 to generate negation arguments for all the selected questions.
-Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting.
-Figure 5 shows a few examples from different categories in the GaslightingBench.
-4 Result Analysis
-Figure 6: Accuracy drop across different categories in GaslightingBench.
 3.1 Performance Comparison
@@
 However, larger models also exhibited more substantial drops after negation, with Qwen2-VL-72B-Instruct seeing a drastic 46.02% decline compared to Qwen2-VL-7B-Instruct’s 32.17% drop.
-This suggests that larger models may be more prone to overfitting or misalignment in adversarial scenarios like negation.
-Figure 7: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct.
-Correct responses are highlighted in green, while incorrect responses are marked in red.
+This suggests that larger models may be more prone to adversarial scenarios like negation.
+Figure 6: Accuracy drop across different categories in GaslightingBench.
 Result in GaslightingBench.
@@
 For example, in the second example in MMMU dataset, LLaVA-1.6-7B generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
-Figure 8: Example shows that different options in negation arguments (e.g., “ten," “one," and “eight") affect MLLMs’ ability to maintain their initial correct response.
+Figure 7: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct.
+Correct responses are highlighted in green, while incorrect responses are marked in red.
 3.2 Discussion: Effect of Negation Argument
@@
 Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
+Figure 8: Example shows that different options in negation arguments (e.g., “ten," “one," and “eight") affect MLLMs’ ability to maintain their initial correct response.
+4 Related Work
+Multimodal Large Language Models.
+Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities.
+The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets.
+Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning.
+Open-source counterparts, such as LLaVA li2024llavanext-strong and Qwen2-VL Qwen2VL-7B , have democratized access to advanced multimodal technologies, fostering innovation within the research community.
+The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions.
+Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct.
+Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
+We leave for future work the exploration of advanced reasoning foundation models, such as OpenAI’s O1 and O3, which are specifically designed to improve multi-step reasoning and tool use across modalities.
+Negation Understanding.
+Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development .
+In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert .
+To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding .
+This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation.
+These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts.
+Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can .
+Recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments.
+Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
+In multimodal learning, negation has received comparatively less attention than in NLP.
+Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision .
+For instance, recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance.
+In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
 5 Conclusion
 This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct.
-While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in their alignment and reasoning mechanisms.
+While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in ensuring the trustworthiness of their outputs.
 Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs.

### v4

@@
-Calling a Spade a Heart: Gaslighting Multimodal Large Language Models via Negation
+Benchmarking Gaslighting Negation Attacks Against Multimodal Large Language Models
 Abstract
 Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements in integrating different modalities, excelling in complex understanding and generation tasks.
-Despite their success, MLLMs remain vulnerable to conversational adversarial inputs, particularly negation arguments.
-This paper systematically evaluates state-of-the-art MLLMs across diverse benchmarks, revealing significant performance drops when negation arguments are introduced to initially correct responses.
+Despite their success, MLLMs remain vulnerable to conversational adversarial inputs.
+In this paper, we systematically study gaslighting negation attacks: a phenomenon where models, despite initially providing correct answers, are persuaded by user-provided negations to reverse their outputs, often fabricating justifications.
+We conduct extensive evaluations of state-of-the-art MLLMs across diverse benchmarks and observe substantial performance drops when negation is introduced.
 Notably, we introduce the first benchmark GaslightingBench, specifically designed to evaluate the vulnerability of MLLMs to negation arguments.
 GaslightingBench consists of multiple-choice questions curated from existing datasets, along with generated negation prompts across 20 diverse categories.
-Throughout extensive evaluation, we find that proprietary models such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA.
-However, all evaluated MLLMs struggle to maintain logical consistency under negation arguments during conversation.
-Our findings provide critical insights for improving the robustness of MLLMs against negation inputs, contributing to the development of more reliable and trustworthy multimodal AI systems.
+Throughout extensive evaluation, we find that proprietary models such as Gemini-1.5-flash and GPT-4o demonstrate better resilience compared to open-source counterparts like Qwen2-VL and LLaVA, though even advanced reasoning-oriented models like Gemini-2.5-Pro remain susceptible.
+Our category-level analysis further shows that subjective or socially nuanced domains (e.g., Social Relation, Image Emotion) are especially fragile, while more objective domains (e.g., Geography) exhibit relatively smaller but still notable drops.
+Overall, all evaluated MLLMs struggle to maintain logical consistency under gaslighting negation attack.
+These findings highlight a fundamental robustness gap and provide insights for developing more reliable and trustworthy multimodal AI systems.
+Project website: https://yxg1005.github.io/GaslightingNegationAttacks/.
 1 Introduction
-Figure 1: Examples demonstrate that GPT-4o initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments.
-GPT-4o shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting".
-Note: "Gaslighting is the manipulation of someone into questioning their own perception of reality." - Wikipedia.
-Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs yin2023survey ; liu2024visual ; bai2023qwen ; hurst2024gpt .
-These models leverage cutting-edge advancements in Large Language Models (LLM) zhao2023survey and multimodal learning radford2021learning ; li2023blip ; liu2024improved , enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning.
-Recent breakthroughs, such as the introduction of proprietary models GPT-4o hurst2024gpt and Claude-3.5-Sonnet anthropic2024claude , have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning.
-Additionally, open-source models like DeepSeek-V3, DeepSeek-R1 liu2024deepseek ; guo2025deepseek and Qwen series bai2023qwen ; qwen2.5 are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs.
+Multimodal Large Language Models (MLLMs) have achieved remarkable progress in understanding and generating language grounded in multimodal contexts, such as visual and textual inputs Yin et al.
+(2023); Liu et al.
+(2024d); Bai et al.
+(2023); Hurst et al.
+(2024).
+These models leverage cutting-edge advancements in Large Language Models (LLM) Zhao et al.
+(2023) and multimodal learning Radford et al.
+(2021); Li et al.
+(2023a); Liu et al.
+(2024b), enabling them to excel in diverse tasks, such as image understanding, visual question answering, and multimodal reasoning.
+Recent breakthroughs, such as the introduction of proprietary models GPT-4o Hurst et al.
+(2024) and Claude-Sonnet Anthropic (2024), have demonstrated exceptional performance, establishing new standards in multimodal understanding and reasoning.
+Additionally, open-source models like DeepSeek-V3, DeepSeek-R1 Liu et al.
+(2024a); Guo et al.
+(2025) and Qwen series Bai et al.
+(2023); Yang et al.
+(2024; 2025) are closing the performance gap with proprietary models, reflecting the rapid and competitive evolution of MLLMs.
 These advancements not only highlight the growing capabilities of LLMs, but also suggest that the development trajectory is edging closer to achieving Artificial General Intelligence (AGI).
+Figure 1: Examples demonstrate that GPT-5 initially provides correct answers but incorrectly revises its responses when confronted with user-provided negation arguments.
+GPT-5 shows a tendency to accept misleading inputs, often generating hallucinated explanations to justify the revised answers, a behavior that can be described as a form of “gaslighting”.
+Note: ”Gaslighting is the manipulation of someone into questioning their own perception of reality.” - Wikipedia.
 Despite the impressive advancements, MLLMs exhibit significant vulnerabilities when navigating complex conversational challenges, particularly those involving adversarial negation.
-This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers wang2023can ; zhao2025aligning .
-As illustrated in Figure 1, GPT-4o initially provides correct answers for all examples.
-However, when presented with negation arguments by the user, GPT-4o often accepts these arguments, revising its answers incorrectly.
-Moreover, it even generates hallucinated explanations to justify these erroneous responses.
-Such susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms sharmatowards .
+This issue becomes evident when models struggle to critically analyze and resist unfaithful arguments, resulting in erroneous reversals of their initially correct answers Wang et al.
+(2023a); Zhao et al.
+(2025).
+As illustrated in Figure 1, GPT-5 initially provides correct answers for all examples.
+However, when presented with negation arguments by the user, GPT-5 often accepts these arguments, revising its answers incorrectly.
+Moreover, it even generates hallucinated explanations to justify these erroneous responses, which we define as gaslighting negation attacks.
+Importantly, such attacks may be injected either intentionally (e.g., malicious manipulation) or unintentionally (e.g., casual user disagreement), making them a pervasive risk in real-world interactions.
+This susceptibility not only undermines their reliability but also exposes fundamental weaknesses in their reasoning and alignment mechanisms Sharma et al.
+(2024).
 These limitations are particularly concerning in high-stakes applications like healthcare diagnostics, autonomous decision-making, and content moderation, where the ability to maintain logical consistency and resist manipulation is crucial.
-Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models huang2024trustllm ; liu2023trustworthy .
-Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after negation, reported as average accuracy across eight benchmarks-MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench , PoPE li2023evaluating-Pope , ChartQA masry2022chartqa , AI2Diagram kembhavi2016diagram and MathVista lu2023mathvista ).
+Addressing these challenges requires a deeper understanding of how MLLMs process and align multimodal inputs, paving the way for more reliable and trustworthy models Huang et al.
+(2024); Liu et al.
+(2023a).
+To comprehensively understand the limitations of MLLMs in handling gaslighting negation attacks, this paper presents the first systematic study by conducting extensive evaluations across eight multimodal benchmarks and state-of-the-art models.
+The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU Yue et al.
+(2024a) and MMBench Liu et al.
+(2025), chart dataset such as ChartQA Masry et al.
+(2022), to math dataset MathVista Lu et al.
+(2023), providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation.
+The proprietary models, such as Gemini, GPT-4o and Claude-Sonnet, alongside open-source counterparts like Qwen-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy.
+As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to these negation attacks, with varying degrees of impact observed across different datasets and MLLMs.
+Figure 2: Comparison of MLLMs’s performance before (i.e., initial answers) and after gaslighting negation attack, reported as average accuracy across eight benchmarks-MME Fu et al.
+(2023), MMMU Yue et al.
+(2024a), MMMUPro Yue et al.
+(2024b), MMBench Liu et al.
+(2025), PoPE Li et al.
+(2023b), ChartQA Masry et al.
+(2022), AI2Diagram Kembhavi et al.
+(2016) and MathVista Lu et al.
+(2023)).
 The results highlight the substantial accuracy drop across all models when negation is introduced.
-The access dates for GPT-4o and Claude-3.5-Sonnet are from 3-March-2025 to 7-March-2025, while results for Gemini-1.5-flash are based on access from 13-Jan-2025 to 18-Jan-2025.
 More detailed results are available in Table 1.
-To comprehensively understand the limitations of MLLMs in handling negation, this paper presents the first systematic study by conducting extensive evaluations across eight multimodal benchmarks and state-of-the-art models.
-The benchmarks cover diverse datasets, ranging from general multimodal datasets such as MMMU yue2023mmmu and MMBench liu2025mmbench , chart dataset such as ChartQA masry2022chartqa , to math dataset MathVista lu2023mathvista , providing a robust framework for assessing the performance of MLLMs under negation arguments in conversation.
-The proprietary models, such as Gemini-1.5-flash, GPT-4o and Claude-3.5-Sonnet, alongside open-source counterparts like Qwen2-VL and LLaVA, allowed for a comparative analysis of reliability and accuracy.
-As shown in Figure 2, through the extensive evaluation, we highlight the widespread vulnerability of MLLMs to negation arguments, with varying degrees of impact observed across different datasets and MLLMs.
-As existing benchmarks primarily evaluate factual accuracy and multimodal reasoning, they fail to systematically assess MLLMs’ susceptibility to negation arguments.
-To address this gap, we introduce GaslightingBench, the first multimodal benchmark designed to evaluate models’ ability to resist misleading negation arguments while maintaining logical consistency.
+As existing benchmarks primarily evaluate factual accuracy and multimodal reasoning, they fail to systematically assess MLLMs’ susceptibility to gaslighting-style manipulations.
+To address this gap, we introduce GaslightingBench, the first multimodal benchmark designed to evaluate models’ ability to resist negation-based adversarial attacks while maintaining logical consistency.
 We curate representative multiple-choice questions (MCQs) from established datasets and generated corresponding negation prompts, resulting in a collection spanning 20 categories and 1,287 samples.
@@
 Notably, while larger models such as Qwen2-VL-72B-Instruct demonstrate higher accuracy in initial responses before negation, it also exhibit significant performance drops after negation, indicating a lack of robustness in handling negation inputs.
-On the contrary, proprietary models such as Gemini-1.5-flash, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments.
-Furthermore, different categories demonstrate varying levels of vulnerability to negation, with “Social Relation" suffering the most significant performance declines, whereas “Geography" experiences comparatively smaller yet still notable drops in our GaslightingBench.
+On the contrary, proprietary models such as Gemini-1.5-flash and reasoning-oriented Gemini-2.5-Pro, show better robustness compared to open-source models, but even they struggle to consistently defend correct answers against misleading negation arguments.
+Furthermore, different categories demonstrate varying levels of vulnerability to negation.
+Degradation is especially severe in subjective or socially nuanced categories (e.g., Social Relation, Image Emotion), while more objective ones (e.g., Geography) are less affected.
+We conjecture that this vulnerability stems from over-alignment with human feedback, which biases models toward user agreement and extends the sycophancy effect observed in LLMs Sharma et al.
+(2024).
 These findings emphasize the urgent need for enhanced training techniques and robust alignment mechanisms to improve the reliability and integrity of MLLMs in real-world applications.
-2 Evaluating Multimodal Large Language Model via Negation
+2 Related Work
+Multimodal Large Language Models.
+Multimodal Large Language Models (MLLMs) Yin et al.
+(2023) represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities.
+The introduction of contrastive learning techniques in vision-language models such as CLIP Radford et al.
+(2021) enables cross-modal understanding through training on extensive datasets.
+Building upon these foundations, modern MLLMs like GPT-4o Hurst et al.
+(2024), Claude-3.5-Sonnet Anthropic (2024), and Gemini-1.5-flash Team et al.
+(2024) have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning.
+Open-source counterparts, such as LLaVA Li et al.
+(2024) and Qwen2-VL Wang et al.
+(2024), have democratized access to advanced multimodal technologies, fostering innovation within the research community.
+The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions.
+Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct.
+Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
+Negation Understanding.
+Negation, defined as the contradiction or denial of something, is a fundamental aspect of language Croft (1991); Pea (1978).
+Early studies showed that models like BERT struggle with distinguishing affirmation and negation Kassner & Schütze (2019); Ettinger (2020), leading to methods such as unlikelihood training to improve performance Hosseini et al.
+(2021).
+More recent work demonstrates that large language models (LLMs), including GPT-3 and InstructGPT, still fail to reliably recognize and reason over negation Truong et al.
+(2023), often reversing correct beliefs when confronted with invalid arguments Wang et al.
+(2023a).
+Approaches such as bilateral confidence estimation and direct preference optimization Zhao et al.
+(2025) attempt to mitigate this by enforcing faithful consistency against opposing claims.
+In contrast, negation has been less explored in multimodal learning.
+Studies on CLIP-like vision–language models (VLMs) Yuksekgonul et al.
+(2022); Singh et al.
+(2024); Wang et al.
+(2023b); Alhamoud et al.
+(2025) reveal similar weaknesses, with recent work showing that fine-tuning on negation-focused datasets improves performance Alhamoud et al.
+(2025).
+Building on these insights, our work shifts focus to MLLMs in conversational settings, examining how they become misled by unfaithful negation even when their initial answers are correct.
+LLM Attacks.
+A large body of work has studied textual manipulations such as jailbreaks Wei et al.
+(2023); Niu et al.
+(2024), prompt injection Liu et al.
+(2023b) and dark patterns Kran et al.
+(2025).
+Jailbreak attacks are primarily designed to bypass safety constraints, coercing models into producing restricted content (e.g., toxic or unsafe outputs), while prompt injection usually aims to override system instructions or insert malicious goals into the input.
+By contrast, our work focuses on gaslighting negation attacks, which are subtler.
+They do not override task instructions or seek restricted outputs, but instead exploit alignment biases to induce models to reverse correct answers and fabricate justifications.
+This distinction highlights gaslighting negation as a complementary and underexplored failure mode that affects reasoning reliability rather than safety guardrails.
+3 Evaluating Gaslighting Negation Attack for MLLMs
 3.1 Experimental Setup
@@
 These models were chosen to reflect a diverse set of training approaches and capabilities in the multimodal AI landscape.
-Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 .
-These models represent the most powerful multimodal AI capabilities.
+Proprietary Models: Our evaluation included leading proprietary models such as GPT-4o Hurst et al.
+(2024), Claude-3.5-Sonnet Anthropic (2024), and Gemini-1.5-flash Team et al.
+(2024).
+These models represent the leading multimodal AI capabilities.
+Our primary analysis focuses on non-reasoning models, however, we also include Gemini-2.5-Pro in its thinking mode to assess how state-of-the-art reasoning-oriented models perform for the gaslighting negation attack.
 All evaluations for proprietary models are performed via their publicly available APIs.
-Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B liu2024llavanext-llava1.6 , Qwen2-VL-7B-Instruct Qwen2VL-7B , LLaVA-NeXT-8B li2024llavanext-strong and Qwen2-VL-72B-Instruct Qwen2VL-7B .
+Open-source Models: In addition to proprietary models, we included open-source models such as LLaVA-1.6-7B Liu et al.
+(2024c), Qwen2-VL-7B-Instruct Wang et al.
+(2024), LLaVA-NeXT-8B Li et al.
+(2024) and Qwen2-VL-72B-Instruct Wang et al.
+(2024).
 These open-source models showcase the rapid progress in publicly available multimodal AI systems.
@@
 Benchmarks.
-To comprehensively assess the robustness of MLLMs in handling negation, we employ a diverse set of benchmark datasets, including general multimodal datasets MME fu2023mme , MMMU yue2023mmmu , MMMUPro yue2024mmmu-MMMUPro , MMBench liu2025mmbench and PoPE li2023evaluating-Pope , Chart dataset ChartQA masry2022chartqa , diagram dataset AI2Diagram kembhavi2016diagram as well as Math dataset MathVista lu2023mathvista .
-Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to negation.
+To comprehensively assess the robustness of MLLMs in handling gaslighting negation attack, we employ a diverse set of benchmark datasets, including general multimodal datasets MME Fu et al.
+(2023), MMMU Yue et al.
+(2024a), MMMUPro Yue et al.
+(2024b), MMBench Liu et al.
+(2025) and PoPE Li et al.
+(2023b), Chart dataset ChartQA Masry et al.
+(2022), diagram dataset AI2Diagram Kembhavi et al.
+(2016) as well as Math dataset MathVista Lu et al.
+(2023).
+Figure 3: Evaluation pipeline for assessing the robustness of Multimodal Large Language Models (MLLMs) to gaslighting negation attack.
 The pipeline consists of three key stages: (1) Inputs and Initial Answers: MLLMs receive a variety of question formats as input, including Yes/No, Multiple-Choice, and Free-Form, and their initial answers are recorded.
@@
 3.2 Evaluation Pipeline
-As shown in Figure 3, we introduce a structured evaluation pipeline to systematically assess MLLMs’ vulnerability to negation by measuring their performance before and after exposure to negation arguments.
+As shown in Figure 3, we introduce a structured evaluation pipeline to systematically assess MLLMs’ vulnerability to gaslighting negation attacks by measuring their performance before and after exposure to negation arguments.
 The process consists of three key steps: First, questions and corresponding options (if applicable) from the original datasets are provided to the MLLMs, and their initial responses are recorded.
@@
 Negation arguments are generated by rewriting the original question into an opposite statement based on the answer.
-If the answer is “Yes.", the question is rewritten as a negative statement and vice versa.
-For example, given the question “Is this photo of Grimspound?" and the answer “Yes.", the negation argument is “This it not a photo of Grimspound." (2) Multiple-Choice Questions.
+If the answer is “Yes.”, the question is rewritten as a negative statement and vice versa.
+For example, given the question “Is this photo of Grimspound?” and the answer “Yes.”, the negation argument is “This it not a photo of Grimspound.” (2) Multiple-Choice Questions.
 For multiple-choice questions, negation arguments are produced by randomly selecting an incorrect option, and presenting it in a negated or contradicted form with the correct answer.
@@
 left D.
-the camera", as well as the answer “D".
-The negation argument could be “The cat is facing right." (option A).
+the camera”, as well as the answer “D”.
+The negation argument could be “The cat is facing right.” (option A).
 (3) Free-Form Questions.
@@
 For instance, give the question “What is the maximum value of y?
-" and answer “5", the negation argument is “The maximum value of y is not 5."
+” and answer “5”, the negation argument is “The maximum value of y is not 5.”
 Post-processing.
@@
 Figure 4: The category distribution of GaslightingBench with 20 categories and 1,287 samples.
-Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to negation arguments.
-Figure 5: Examples from different categories in the GaslightingBench. The green-highlighted option is correct, while a randomly chosen incorrect option is used to generate the negation argument.
+Each category is carefully curated from existing datasets to ensure balanced representation and broad coverage, providing a comprehensive evaluation dataset for assessing MLLMs’ vulnerabilities to gaslighting negation attacks.
 3.3 Gaslighting Benchmark Construction
-To comprehensively evaluate the impact of negation on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments.
+To comprehensively evaluate the impact of gaslighting negation attacks on MLLMs, we introduce the first gaslighting benchmark, GaslightingBench, designed to ensure broad coverage, category balance, and the ability to expose MLLMs’ vulnerabilities to negation arguments.
 GaslightingBench exclusively uses Multiple-Choice Questions (MCQ) to facilitate structured evaluation, as MLLMs generally provide more consistent responses in MCQs than in open-ended formats, while MCQs remain more complex than binary Yes/No questions.
@@
 (2) To maintain a sufficient number of samples per category, we manually merge semantically similar categories.
-For example, in MMBench,“Attibute_comparison" and “Attribute_recognition" are combined under “Attribute" to ensure broader coverage.
+For example, in MMBench,“Attibute_comparison” and “Attribute_recognition” are combined under “Attribute” to ensure broader coverage.
 (3) For categories present in multiple datasets, we incorporate samples from different sources to ensure representativeness.
-For instance, our “Math" category consists of samples from both AI4MATH and MMMU-Math.
+For instance, our “Math” category consists of samples from both AI4MATH and MMMU-Math.
 Second, we employ the same approach detailed in section 3.2 to generate negation arguments for all the selected questions.
+In addition, we retain only questions that are answerable and visually grounded, support faithful negation prompts, and contain high-quality distractors.
+Categories are further consolidated via a manual ontology-driven process to reduce sparsity and fragmentation.
 Ultimately, as shown in Figure 4, GaslightingBench comprises 20 categories and 1,287 samples, covering a wide range of topics, such as Business, Medicine, Image Emotion, and Counting.
 Figure 5 shows a few examples from different categories in the GaslightingBench.
+Figure 5: Examples from different categories in the GaslightingBench. The green-highlighted option is correct, while a randomly chosen incorrect option is used to generate the negation argument.
 4 Result Analysis
+4.1 Performance Comparison
+Table 1 lists the results across all evaluated benchmarks, comparing the accuracy of selected MLLMs before (i.e., initial answers) and after the introduction of negation arguments.
+Overall, all the MLLMs exhibit significant performance declines even though the accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts.
+The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60% (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
 Model Negation Dataset average
-MME fu2023mme MMMU yue2023mmmu MMMUPro yue2024mmmu-MMMUPro AI2Diagram kembhavi2016diagram MathVista lu2023mathvista ChartQA masry2022chartqa PoPE li2023evaluating-Pope MMBench liu2025mmbench
+MME MMMU MMMUPro AI2Diagram MathVista ChartQA PoPE MMBench
 Open-Source
@@
 […]
-before 86.02 50.37 27.05 79.83 60.00 75.32 87.94 82.26 68.42
+[…]
 […]
@@
 […]
-before 91.70 60.34 33.64 85.30 67.10 80.60 87.84 79.81 73.29
+[…]
 […]
@@
 […]
-before 69.17 62.07 36.42 75.84 54.60 72.68 85.46 77.69 66.74
+[…]
 […]
@@
 The performance drop is highlighted in red.
-Model LLaVA-1.6-7B liu2024llavanext-llava1.6 Qwen2-VL-7B-Instruct Qwen2VL-7B LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-72B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5 GPT-4o hurst2024gpt Claude-3.5-Sonnet anthropic2024claude
-before negation 59.13 76.85 71.33 77.08 74.13 69.15 78.17
-[…]
-Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after negation. The performance drop is highlighted in red.
-Model LLaVA-1.6-7B liu2024llavanext-llava1.6 LLaVA-NeXT-8B li2024llavanext-strong Qwen2-VL-7B-Instruct Qwen2VL-7B Gemini-1.5-flash team2024gemini1.5
-before negation 31.77 42.98 50.37 57.39
 […]
 […]
-Table 3: Performance comparison of Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset.
-The performance drop compared to before negation is indicated in red.
-3.1 Performance Comparison
-Table 1 lists the results across all evaluated benchmarks, comparing the accuracy of selected MLLMs before (i.e., initial answers) and after the introduction of negation arguments.
-Overall, all the MLLMs exhibit significant performance declines even though the accuracy before negation is relatively high, with open-source models generally showing greater vulnerability compared to proprietary counterparts.
-The average accuracy drop ranges from 9.62% (LLaVa-1.6-7B in MMMUPro dataset) to 66.60% (Qwen2-VL-72B-Instruct in MMBench dataset), highlighting the widespread challenges MLLMs encounter when handling negation-based inputs in conversation.
+[…]
+Table 2: Results of MLLMs in our GaslightingBench, comparing each model’s performance before (i.e., initial answers) and after gaslighting negation attack.
+Gemini-2.5-Pro (highlighted in gray) is evaluated in thinking mode with multi-step reasoning, whereas all other models are non-reasoning.
+The performance drop is highlighted in red.
 Open-Source vs.
@@
 In addition, proprietary models like GPT-4o do not consistently outperform open-source models across all datasets.
-On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown." For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as "unknown." Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
+On tasks involving true/false or multiple-choice questions, such as those in MME and AI2Diagram, GPT-4o often fails to provide a definitive answer, frequently responding with “unknown.” For instance, in MME, GPT-4o incorrectly answered nearly 150 out of 400 questions related to artwork, many of which were labeled as ”unknown.” Furthermore, GPT-4o struggles with datasets like MathVista and ChartQA, where large open-source models like Qwen2-VL-72B outperformed it.
 Comparison between Larger and Smaller Models.
@@
 Table 2 presents the evaluation results of MLLMs on GaslightingBench, highlighting their accuracy before and after the introduction of negation arguments.
-Similarly, across all models, significant performance degradation is observed, reaffirming the susceptibility of MLLMs to negation.
+Similarly, across all models, significant performance degradation is observed, reaffirming the susceptibility of MLLMs to gaslighting negation attack.
+While our main analysis focuses on non-reasoning models, we also evaluate an advanced reasoning-oriented MLLM to test the generality of this vulnerability.
+Specifically, we examine Gemini-2.5-Pro, currently first-ranked MLLM on the LMArena leaderboard  [^1] https://lmarena.ai/leaderboard/vision and equipped with a dedicated “thinking” mode for multi-step reasoning.
+Gemini-2.5-Pro achieves strong performance before negation (87.74) and shows relatively smaller degradation compared to the other non-reasoning models.
+However, despite its chain-of-thought reasoning, the model still exhibits notable vulnerability under gaslighting negation, with accuracy dropping to 70.86.
+These results reinforce our key finding: even the most capable MLLMs struggle to maintain consistent reasoning in the face of gaslighting negation, highlighting a fundamental limitation across models.
 As shown in Figure 6, different categories show varying degrees of susceptibility to negation arguments.
-Notably, “Social Relation" and “Image Emotion" experience the most severe drops, with accuracy declining by over 50%, indicating their high sensitivity to negation arguments.
-In contrast, “Geography" and “Chemistry" demonstrate comparatively lower declines but still suffer more than a 25% accuracy reduction, emphasizing that no category remains unaffected.
-Similar observations also appear in other existing datasets, for instance, MMMU dataset, which contains 30 subjects such as “accounting" and “Art”, the largest performance drop is observed in “Art Theory", while “Music" experiences the smallest decline.
-Similarly, in the MME dataset, “Counting" shows the most significant drop, whereas “Numerical Calculation" is the least affected.
+Notably, “Social Relation” and “Image Emotion” experience the most severe drops, with accuracy declining by over 50%, indicating their high sensitivity to negation arguments.
+In contrast, “Geography” and “Chemistry” demonstrate comparatively lower declines but still suffer more than a 25% accuracy reduction, emphasizing that no category remains unaffected.
+Similar observations also appear in other existing datasets, for instance, MMMU dataset, which contains 30 subjects such as “accounting” and “Art”, the largest performance drop is observed in “Art Theory”, while “Music” experiences the smallest decline.
+Similarly, in the MME dataset, “Counting” shows the most significant drop, whereas “Numerical Calculation” is the least affected.
 These results highlight the uneven but widespread impact of negation arguments across different domains.
+Figure 7: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct.
+Correct responses are highlighted in green, while incorrect responses are marked in red.
 Qualitative Analysis.
@@
 For example, in the second example in MMMU dataset, LLaVA-1.6-7B generates explanations following negation arguments lack grounding in the visual content, highlighting the models’ overconfidence in producing unverified reasoning.
-Figure 7: Qualitative examples illustrating how various MLLMs respond to negation arguments when their initial answers are correct.
-Correct responses are highlighted in green, while incorrect responses are marked in red.
-3.2 Discussion: Effect of Negation Argument
+4.2 Discussion
+Effect of negation types.
+Negation is linguistically diverse and extends beyond the neutral forms introduced in Section 3.2.
+To examine this, we incorporated two additional variants into GaslightingBench: (i) anger-style negation, where the user conveys emotionally charged disbelief (e.g., “I can’t believe you made such a basic mistake!”), and (ii) authority-style negation, where the user appeals to an external authority (e.g., “The professor said your answer is incorrect.”).
+As shown in Table 3, we observed a larger performance decline compared to neutral negation.
+This indicates that emotionally or authoritatively framed challenges can further erode model reliability, likely by amplifying the model’s deference to perceived user authority.
+Model LLaVA-1.6-7B Qwen2-VL-7B LLaVA-NeXT-8B Qwen2-VL-72B Gemini-1.5-flash GPT-4o Claude-3.5-Sonnet
+Before negation 59.13 76.85 71.33 77.08 74.13 69.15 78.17
+[…]
+[…]
+[…]
+Table 3: Performance of MLLMs on GaslightingBench under different negation types.
+Model confidence under gaslighting negation attacks.
+To better understand the internal behavior of MLLMs under gaslighting negation attacks, as shown in Table 4, we conduct a confidence-based analysis using model-reported probability scores for Gemini-1.5-flash and Qwen-2-VL-7B on both GaslightingBench and MMMU.
+We group predictions by their correctness and whether they occurred before or after negation, then compute the average confidence scores.
+On the one hand, for both models, confidence scores are generally higher for correct predictions than for incorrect ones, especially after negation, indicating some degree of internal calibration.
+On the other hand, we observe confidence drop in incorrect answers after negation, particularly in Qwen-2-VL-7B (from 90.6 to 74.4 on GaslightingBench), suggesting the model becomes less confident when misled.
+However, confidence in incorrect responses remains relatively high, especially in Gemini-1.5-flash, which maintains 91.9 average confidence even when wrong after negation, indicating a risk of confident hallucination.
+These findings indicate that while models may exhibit partial uncertainty when manipulated, they can still produce incorrect yet high-confidence outputs.
+This suggests the need for calibrated uncertainty modeling in MLLMs under adversarial dialogue settings.
+Model Gemini-1.5-flash Qwen-2-VL-7B
+Negation Before After Before After
+correct incorrect correct incorrect correct incorrect correct incorrect
+GaslightingBench 95.9 91.5 92.9 87.2 90.2 90.6 81.2 74.4
+MMMU 93.6 91.0 94.8 91.9 90.5 91.0 84.5 87.2
+Table 4: Average confidence scores of Gemini-1.5-flash and Qwen2-VL-7B on GaslightingBench and MMMU, grouped by correctness before and after gaslighting negation.
+Why are MLLMs prone to gaslighting negation attacks?
+Our results suggest that the vulnerability of MLLMs to gaslighting-style negation stems primarily from over-alignment with human feedback in multimodal reasoning tasks.
+Many state-of-the-art MLLMs are trained using human preference optimization techniques, such as instruction tuning and reinforcement learning from human feedback (RLHF).
+While these techniques improve model helpfulness and cooperation, they also introduce a bias toward agreeing with user input, especially in conversational contexts.
+This can result in over-deference, where models revise initially correct answers simply in response to user disagreement, regardless of the factual correctness of the user’s claim.
+This behavior parallels the phenomenon of sycophancy observed in LLMs Sharma et al.
+(2024) and appears to extend into the multimodal setting as well.
+Our category-level analysis in Figure 6 further shows that the performance degradation is especially pronounced in subjective or socially nuanced categories (e.g., “Social Relation”, “Image Emotion”), whereas more objective domains like “Geography” show comparatively smaller drops.
+This suggests that in addition to alignment bias, task uncertainty and subjective ambiguity also play a role in model susceptibility.
+Our primary goal in this paper is to establish a rigorous and extensible framework for evaluating and benchmarking MLLM robustness under gaslighting negation attacks.
+Mitigating this issue will likely require more fine-grained alignment strategies, such as distinguishing faithful correction from invalid contradiction, and improving calibration of confidence in multimodal predictions Zhao et al.
+(2025).
+We leave these directions for future work.
+5 Conclusion
+This paper has revealed critical vulnerabilities in MLLMs when exposed to gaslighting negation attacks, where correct answers are overturned by misleading prompts.
+Our comprehensive evaluation across diverse benchmarks demonstrates that this susceptibility is widespread, affecting both proprietary and open-source models, including those with advanced reasoning capabilities.
+These findings highlight a fundamental gap in model robustness and raise important concerns for the trustworthiness of MLLMs in real-world applications.
+By introducing GaslightingBench, we provide the first systematic benchmark for evaluating this vulnerability, offering a foundation for future work on robustness and alignment.
+We encourage further exploration of fine-grained training strategies, confidence calibration, and evaluation frameworks that can distinguish valid corrections from misleading contradictions.
+Addressing these challenges is essential for advancing multimodal AI toward systems that maintain logical consistency and reliability under adversarial conditions.
+Reproducibility Statement
+We will release the full code, evaluation pipeline, negations for all of datasets we used, and GaslightingBench benchmark.
+Proprietary models were accessed via public APIs, and open-source models were evaluated with fixed checkpoints.
+The prompt templates are provided in the Appendix.
+All datasets evaluated in this paper are publicly available.
+Appendix A Appendix
+A.1 Performance Comparison Across Question Types
+To investigate how question formats affect MLLMs’ robustness to gaslighting negation attack, we conduct a format-level analysis using the MathVista dataset, which contains Yes/No, Multiple-Choice (MCQ), and Open-Ended question types.
+We evaluate two models: Qwen2-VL-7B and LLaVA-NEXT-8B.
+As shown in the Table 5, based on the performance before negation, the relative difficulty of question formats follows the order: Open-Ended >> MCQ >> Yes/No, which aligns with expectations given that open-ended questions involve a broader and less constrained answer space.
+Nevertheless, open-ended questions demonstrate greater resilience to negation compared to Yes/No and MCQ formats.
+For example, Qwen2-VL-7B maintains 37.61% accuracy on open-ended questions after negation, compared to 26.72% for MCQs and 40.11% for Yes/No.
+Model Qwen2-VL-7B LLaVA-NEXT-8B
+Question Type Open-Ended MCQs Yes/No Open-Ended MCQs Yes/No
+Before Negation 50.87 57.02 89.83 25.43 34.16 61.02
+After Negation 37.61 26.72 40.11 14.13 9.09 19.21
+Table 5: Accuracy of Qwen2-VL-7B and LLaVA-NEXT-8B on the MathVista dataset across different question formats (Open-Ended, MCQ, Yes/No) before and after gaslighting negation.
+A.2 Impact of LLM Choice on Negation Generation
+Although Llama3-8B-Instruct is not among the strongest models overall, we adopt it for negation prompt generation (Section 3.2) because the task is relatively constrained: the model is given the original question, answer, and candidate options (if applicable) and asked to produce a logically negated statement.
+In this controlled setting, even moderately sized LLMs can generate reliable outputs.
+To validate quality, we regenerated all negation prompts using a flagship model, Gemini-2.5-Pro, and compared them against those from Llama3-8B-Instruct.
+Using Qwen3 embeddings to measure semantic similarity, we obtained an average score of 0.90, confirming that the two sets of prompts are highly consistent, with negligible differences in quality.
+A.3 Scope of GaslightingBench’s Question Design
+The choice to base GaslightingBench primarily on Multiple-Choice Questions (MCQs) was deliberate, as discussed in Section 3.3.
+MCQs offer a balance of semantic complexity, evaluation consistency, and annotation reliability.
+They also enable precise measurement of model behavior under negation with minimal ambiguity, making them a practical foundation for probing manipulations such as gaslighting.
+That said, assessing negation vulnerability across diverse question formats is equally important.
+Our evaluation pipeline (Section 3.2) already supports Yes/No and Free-Form questions, which we evaluate using model confidence scores or semantic similarity.
+Indeed, the results in Table 1 span all eight datasets, covering a mix of MCQs, binary (Yes/No), and open-ended formats (e.g., MME and MathVista).
+While the initial release of GaslightingBench focuses on MCQs, we will also provide a comprehensive set of negation prompts for all eight evaluated datasets, including non-MCQ formats.
+This will allow the community to extend evaluations to additional question types and investigate how gaslighting negation generalizes beyond MCQs.
+Model LLaVA-1.6-7B LLaVA-NeXT-8B Qwen2-VL-7B-Instruct Gemini-1.5-flash
+before negation 31.77 42.98 50.37 57.39
+[…]
+[…]
+Table 6: Performance comparison of Multimodal Large Language Models (MLLMs) under two distinct negation strategies from the MMMU dataset.
+The performance drop compared to before negation is indicated in red.
+A.4 Effect of Negation Argument
 To investigate the impact of different negation strategies, we conduct experiments using two distinct negation approaches in the MMMU dataset: (1) Direct Negation Using Options.
-This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C." (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C)." As shown in Table 3, all the MLLMs show higher accuracy for the second strategy after negation.
+This strategy refers to directly negating the initial answer by stating the correct option explicitly, such as “The correct answer is C.” (2) Negation with Descriptive Content of Option represents that we negate the initial answer by providing a description of an incorrect option, such as “… is Explicit Themes (the content of option C).” As shown in Table 6, all the MLLMs show higher accuracy for the second strategy after negation.
 The result suggests that providing additional context can help models better retain reasoning integrity when challenged with negation arguments.
 Since descriptive content more effectively demonstrates the capabilities of MLLMs compared to direct options, we adopt the second strategy for all experiments.
+In addition, we also adapt our GaslightingBench to support initial-query negation by injecting the negation statement into the original prompt.
+The results in Table 2 present a comparison on LLaVA-1.6-7B.
+We observe that initial-query negation leads to comparable or slightly worse performance than conversational negation, suggesting that negation vulnerability is pervasive regardless of prompt structure.
+Model Negation GaslightingBench
+LLaVA-1.6-7B Before 59.13
+After(conversational negation) 27.20
+Initial-query negation 21.29
+Table 7: Performance of LLaVA-1.6-7B on GaslightingBench between conversational and initial-query negation.
 Furthermore, as shown in Figure 8, the choice of negation arguments significantly affects MLLMs’ performance.
-For instance, when the negation argument includes “ten" or “eight," only two models maintain the correct response of “four," whereas using “one" as the negation argument allows six models to retain the correct answer.
+For instance, when the negation argument includes “ten” or “eight,” only two models maintain the correct response of “four,” whereas using “one” as the negation argument allows six models to retain the correct answer.
 This highlights the varying sensitivity of MLLMs to different negation formulations.
@@
 Exploring these variations will be our future work to comprehensively assess model vulnerabilities and enhance the robustness of MLLMs under diverse negation conditions.
-Figure 8: Example shows that different options in negation arguments (e.g., “ten," “one," and “eight") affect MLLMs’ ability to maintain their initial correct response.
-4 Related Work
-Multimodal Large Language Models.
-Multimodal Large Language Models (MLLMs) yin2023survey represent a significant evolution in artificial intelligence, integrating multiple data modalities—such as text, images, and audio—to enhance understanding and generation capabilities.
-The introduction of contrastive learning techniques in vision-language models such as CLIP radford2021learning enables cross-modal understanding through training on extensive datasets.
-Building upon these foundations, modern MLLMs like GPT-4o hurst2024gpt , Claude-3.5-Sonnet anthropic2024claude , and Gemini-1.5-flash team2024gemini1.5 have integrated vision encoders with large language models, achieving state-of-the-art performance in tasks including visual question answering and image reasoning.
-Open-source counterparts, such as LLaVA li2024llavanext-strong and Qwen2-VL Qwen2VL-7B , have democratized access to advanced multimodal technologies, fostering innovation within the research community.
-The typical training process for these models involves two stages: vision-language alignment pretraining, which maps visual features to the language model’s embedding space using large-scale image-text pairs, and visual instruction tuning, which fine-tunes the model to handle diverse visual instructions.
-Despite these advancements, this paper reveals that MLLMs are vulnerable to misleading negation arguments, even when their initial responses are correct.
-Addressing such weaknesses is essential for improving the reliability and robustness of MLLMs in practical applications.
-We leave for future work the exploration of advanced reasoning foundation models, such as OpenAI’s O1 and O3, which are specifically designed to improve multi-step reasoning and tool use across modalities.
-Negation Understanding.
-Negation, defined as the contradiction or denial of something, is a fundamental aspect of language croft1991evolution ; pea1978development .
-In natural language processing (NLP), Pretrained Language Models like BERT have been shown to struggle with distinguishing affirmation and negation cloze questions kassner2019negated ; ettinger2020bert .
-To address this, methods like unlikelihood training have been introduced to improve BERT’s ability to understand negation using negated sentences from existing datasets hosseini2021understanding .
-This line of research has advanced further with studies like truong2023language , which show large language models (LLM), including GPT3 and InstructGPT, also exhibit significant limitations in handling negation.
-These models often fail to recognize the presence of negation, struggle to accurately interpret the lexical semantics of negation, and face challenges in maintaining logical consistency and reasoning effectively when confronted with negated contexts.
-Moreover, LLMs frequently struggle to maintain logical consistency when defending correct beliefs against invalid arguments, raising concerns about their alignment and depth of understanding wang2023can .
-Recent work zhao2025aligning addresses the challenge of maintaining faithful integrity in large language models, which leverage bilateral confidence estimation and direct preference optimization to ensure LLMs remain consistent with correct statements while adapting to valid opposing arguments.
-Building on the insights of wang2023can ; zhao2025aligning , this paper extends the investigation into the impact of opposing arguments in MLLMs.
-In multimodal learning, negation has received comparatively less attention than in NLP.
-Existing works have been explored to improve the ability to understand opposing arguments of CLIP-like vision-language models (VLMs) yuksekgonul2022and ; singh2024learn ; wang2023clipn ; alhamoud2025vision .
-For instance, recent work alhamoud2025vision highlights the limitations of VLMs in understanding negation across tasks such as retrieval and multiple-choice questions with negation, while demonstrating that fine-tuning with negation-focused datasets can significantly improve performance.
-In contrast, this paper focuses on evaluating the negation understanding of MLLMs in conversational settings, particularly when the models’ original answers are correct but become misled by unfaithful negation arguments during conversation.
-5 Conclusion
-This paper has revealed critical vulnerabilities in MLLMs when exposed to negation arguments in conversation when the initial answers are correct.
-While these models demonstrate remarkable advancements in multimodal understanding and reasoning, their susceptibility to misleading prompts highlights fundamental weaknesses in ensuring the trustworthiness of their outputs.
-Through comprehensive evaluation across diverse benchmarks, we emphasize the importance of addressing these challenges to enhance the reliability and robustness of MLLMs.
-This paper also provides valuable insights for future research aimed at developing more trustworthy AI systems capable of maintaining logical consistency under adversarial negation conditions.
-We encourage further exploration of advanced training techniques and evaluation frameworks to mitigate the risks posed by such adversarial inputs, ultimately advancing the state of multimodal AI toward more reliable and trustworthy real-world applications.
+Figure 8: Example shows that different options in negation arguments (e.g., “ten,” “one,” and “eight”) affect MLLMs’ ability to maintain their initial correct response.
+A.5 Prompt for gaslighting negation attack evaluation
+The prompt template for generation negation statements and evaluation prompt templates are presented in Figures 9 and 10, respectively.
+Figure 9: Prompt templates used for generating negation statements across different question formats.
+Figure 10: Evaluation prompt templates used for different MLLMs across question formats.
+A.6 The Use of Large Language Models (LLMs)
+In this paper, LLMs/MLLMs are used in three ways: (i) Negation prompt generation.
+We employ Llama3-8B-Instruct to generate neutral negation prompts, chosen for efficiency and accessibility.
+To ensure quality, we cross-validated with Gemini-2.5-Pro and found high semantic similarity (0.90 via Qwen3 embeddings), confirming consistency across models.
+(ii) Model evaluation.
+Proprietary systems such as GPT-4o, Claude-3.5-Sonnet and Gemini-1.5/2.5 were accessed via official APIs, while open-source baselines such as Qwen-VL and LLaVA were tested on released checkpoints.
+All inference settings, hyperparameters, and evaluation pipelines are documented to support reproducibility.
+(iii) Paper writing refinement.
+We used LLM assistance to polish the writing and improve readability, while all technical content, methodology, and experimental design were authored and validated by the research team.